Bryan's boat travels 15 km/h in still water.He motors 140 km downstream in the same time it takes to travel 35 km upstream. What is the speed of the river?
Answer by richwmiller(17219)

r*t=d
(15-c)*t=35
15*t-c*t=35
(15+c)*t=140
15*t-c*t=35
15*t+c*t=140
add
30t=175
t=175/30
t=5.83333333 5 hr 50 minutes hr
15*5.83333333+c*5.83333333=140
87.5+c*5.83333333=140
c*5.83333333=52.5
c=52.5/5.83333333
c=9.0 mph current
check
(15+c)*t=140
(15+9)*5.83333333=140
(24)*5.83333333=140
140.0=140
(15-c)*t=35
(15-9)*5.83333333=35
(6)*5.83333333=35
35.0=35
ok